OpenStudy (anonymous):

3, __, __, __, __, 96

OpenStudy (anonymous):

arithmetic means or geometric means?

OpenStudy (anonymous):

the function rule: \(\Large f(x)= 3\times 8n\ (0\le n \le 4)\) \(\Large \color{blue}{n=0}\ f(x)=3\) \(\Large \color{blue}{n=1}\ f(x)=3\times 8=24\) \(\Large \color{blue}{n=2}\ f(x)=3\times 8\times 2=48\) \(\Large \color{blue}{n=3}\ f(x)=3\times 8\times 3=72\) \(\Large \color{blue}{n=4}\ f(x)=3\times 8\times4=96\)

OpenStudy (anonymous):

4x + 1 < 2x + 5 HLP PLZ

OpenStudy (anonymous):

\(\Large 4x + 1 < 2x + 5\) \(\Large 4x-2x<5-1\) \(\Large 2x<4\) \(\Large x<2\)